40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Thu Aug 26 01:01:34 2010, Zach, Laser, GYRO, gyro back to gyro configuration 2010-08-25_13.19.11.jpg2010-08-25_13.20.31.jpg
    Reply  Fri Aug 27 11:51:41 2010, Zach, Laser, GYRO, gyro back to gyro configuration 
Message ID: 997     Entry time: Fri Aug 27 11:51:41 2010     In reply to: 986
Author: Zach 
Type: Laser 
Category: GYRO 
Subject: gyro back to gyro configuration 

I got the PLL back up and running yesterday before I left. I spent a disproportionate amount of time not knowing that the loop was not closed because "external source" was not on, despite that FM modulation was on and the source was set to external. I'm surprised there is no "modulation source to modulation ON/OFF" or "Displayed instrument settings to actual instrument settings ON/OFF" toggle on top of this.

The gyro is now in a state where it auto-acquires lock in both directions (though it sometimes catches on the TEM10), and the beat frequency is well within the PLL's lock range, so we get a gyro signal automatically. This signal (external voltage on the VCO) is acquired at GENERIC_DOF6_OUT, while TRANS DC is picked up at GENERIC_DOF7_OUT. I think DOF8_OUT has the actuation signal to the AOM VCO, but I don't remember for sure.

I made changes to various settings along the way. Here are the current ones:

  • CW PDH (AOM): G = 7, inv ON, boost ON, phi = 5. There is now a 20 dB attenuator at the output to the VCO. This seemed to help with keeping the gain at a reasonable value while not limiting our range on the VCO in lowering the deviation.
  • CCW PDH: G = 7, inv OFF, boost ON, phi = 4
  • AOM VCO: fnom = 47.451 MHz, dev = 100 kHz
  • PDH LO: fmod = 18 MHz. I tried using 33 MHz but the signal out of the (low-bandwidth) CW PD was puny and this generated noise downstream. This frequency appears devoid of other-modely interference.
  • PLL: fnom = 94.895719 MHz, dev = 375 kHz, SR560 G = 1 with no filtering (i.e. doing nothing)

I was not able to take a spectrum with DTT as the FB was dead, but I dumped the gyro signal into the SR785 as I was monkeying in the parameter space, and I was able to get the gyro noise at and slightly below 10 Hz to be on par with or ~ a factor of 2 or 3 better than it was before. This is taking into account that the previous calibration was off by a factor of 2*pi (lambdaP/4A takes optical Hz to rad/s, not RPS), so the noise can be made roughly 10-20 times lower than it used to be.

That said, it bothers me that we have no way of knowing at present if we are increasing our SNR or just reducing both noise AND signal whenever we make a change to a setting. I could have "made the noise much lower" by simply reducing the gain in the CW path, but this is not what would really be happening.

My issues with the PLL and settings-tweaking, along with Pinkesh's defense, precluded me from taking the transfer functions I wanted to take. These are the first step in characterizing our loops and ultimately in improving our SNR. They must be done before anything else.

Quote:

 I put the gyro back into the normal configuration today. It is now locking in both directions again. The bandwidth seems slightly improved, as I didn't have to be quite as careful when working on the table to keep it from dropping, but it is far from good, and the 200kHz+ oscillation still shows up with enough gain. I had set up the PLL but was having problems locking it when I had to leave.

A couple changes:

  • I kept the high-speed PDA10A that I used for the linear cavity in place as the CCW REFL PD, instead of the old PDA36A. We eventually want to have the same two PDs for both REFLs, but this doesn't really matter at the moment, and we might as well maximize our optical gain in the main locking loop while using the design modulation frequency of 33 MHz. The CW loop is still a 17-MHz diode, but this loop seems to like the lowest gain setting anyway, and we may end up wanting to reduce lower gain limit of the CW PDH box.
  • The BS1 in the TRANS readout was a 45P, but we are using S, so I changed it to a 45S.

I took some time to label nearly all of the cables we work with at the controls interface. The REFL PD signals, the CW & CCW error signals, the PZT and AOM VCO control signals, the TRANS DC signal, the SWEEP drive from the function generator, and two cables going to dedicated DAQ channels have all been labeled. This way we don't have to tug on cables or follow them across the table to figure out what is what. We should be prudent that we don't change them around without relabeling them appropriately

2010-08-25_13.19.11.jpg2010-08-25_13.20.31.jpg

Tomorrow morning I'm going to figure out what's wrong with the PLL, then lock it and take a spectrum to see if things have gotten better with the new PD, new modulation frequency, and fresh realignment. Then I plan to take a bunch of transfer functions to try and diagnose our loop problem. Specifically:

  1. The closed-loop TF of each direction
  2. The OLTF from the SWEEP input of each PDH box to its output

(2) is necessary to obtain the OLTF of the loops from (1), as we will be using the SWEEP input for the excitation, and the TF from SWEEP to OUT is not the same as that from IN to OUT.

 

ELOG V3.1.3-